OpenStudy (anonymous):

what is the solution to the following system? -3x-2y=-12 9x+6y=-9 choices A.(2,1) B.no solution C.(-2,-1) D.infinitely many solutions

OpenStudy (zpupster):

go ahead kadin use what learned earlier, pick substitution or elimination

OpenStudy (anonymous):

pick eihter one?

OpenStudy (zpupster):

sure they both work it just asks for a solution

OpenStudy (anonymous):

okay ill go with elimination

OpenStudy (zpupster):

good choice

OpenStudy (zpupster):

what can you mult the first eqaution by to get rid of x

OpenStudy (anonymous):

by six right? because that would give me 18 and 9 times 2 is 18?

OpenStudy (zpupster):

-3x - 2y = -12 9x + 6y = -9

OpenStudy (zpupster):

i was thinking 3 that would give -9 and 9 when added they cancel

OpenStudy (anonymous):

okay that makes more sense

OpenStudy (anonymous):

-9x - 2y = -12 9x + 6y = -9

OpenStudy (zpupster):

no you have to mult every term by 3 also -2y and -12

OpenStudy (anonymous):

oh so -9x-6y=36

OpenStudy (zpupster):

yes now add the equations together

OpenStudy (anonymous):

do i do it by like terms?

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

0x 0y = 27

OpenStudy (zpupster):

yes it is safe to say no solutions can be found

OpenStudy (anonymous):

okay im starting to understand these much better would you like to help me with another one just to be sure?

OpenStudy (zpupster):

sorry lunch time, you can message i may be bck later today. remember to do your checks if the checks do not work you did something wrong.

OpenStudy (anonymous):

alright thanks!!